Fermi National Laboratory
 
Operation Summary for February 24, 2012 - February 27, 2012
 

NuMI reported receiving 7.92E18 protons on target for the period from 2/20/12 to 2/27/12.

Friday February 24

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, a Booster person began a study at 9:21 AM. Operators reset three MI vacuum pumps (604E and 605A & B) that had tripped off at 9:21 AM. Safety personnel accessed Switchyard (SY) enclosures C, D, and E at 9:29 AM for safety inspections. A Controls person rebooted the SY and Meson frontends at 9:59 AM. Operators searched and secured the Meson MTest areas and Operations established beam to the EDIT School at 12:11 PM. Neutron Therapy Facility personnel reported at 12:19 PM that they had completed treating their last patient for the day. The Booster study ended at 12:19 PM. A Linac Klystron RF station (KRF6) tripped off at 2:18 PM; it reset. At 2:34 PM, the NuMI beam permit tripped off on an incomplete RF indication; it reset. A Booster person began a study at 2:57 PM. At 3:44 PM, a Booster bulk power supply (CEP400) tripped off; Booster personnel investigated and brought CEP400 back online, but a quadrupole magnet power supply (Q803), which had also tripped, would not come back on. A Power Supply person investigated. Operators reported that there were MI studies throughout the afternoon.

During the evening shift, Booster resumed sending beam at 4:42 PM; Power Supply personnel replaced the Q803 power supply. MI studies ended at 5:16 PM. Operations disabled beam to the EDIT School at 6 PM, which completed the school’s run. At 7:40 PM, the NuMI beam permit tripped off due to an incomplete RF gap, but no MI RF station had tripped off; the beam permit reset. At 10:58 PM, a Booster RF station (BRF19) tripped off for the first of three times in about 40 minutes; Operators reset the station. At 11:11 PM, BRF15 tripped off; it reset.

Saturday February 25

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 12:42 AM, two Linac quadrupole magnet power supplies (L:QPS105 and 108) tripped off; they reset. LRF3 tripped off at 1:30 AM; it reset. At 2:52 AM, Operators reported that the Booster RF system as a whole began sparking a lot and BRF19 began tripping off repeatedly; Operators turned off BRF19, but it didn’t help. Operations reduced the amount of beam in the Booster and called in a Booster person. The Booster person asked Operators to turn off one of the fans in the low level RF room since its temperature had dropped. By 4:01 AM, the temperatures had stabilized and Operators had brought the Booster back up to normal beam intensities. The Booster RF system resumed sparking at 4:42 AM. At 6:38 AM, LRF4 tripped off for the first of three times in about 20 minutes; after the last trip it would not reset. Operators called in Linac personnel. Operations put the accelerators into an energy conservation timeline at 7:46 AM.

During the day shift, Linac resumed sending beam around 8:10 AM; Linac personnel found that a fuse wire had popped off. At 8:42 AM, LRF1 tripped off; it reset. Operators completed a 750 KeV line tune at 9:17 AM. Operators loaded protons into the Recycler at 9:25 AM. At 10:20 AM, BRF16 tripped off for the first of two times in 14 minutes; it reset. Operators completed a Booster injection vertical angle bump to improve beam positions at 11:56 AM. At 2:03 PM, LRF3 tripped off; it reset.

During the evening shift, at 5:21 PM, KRF1 tripped off; it reset. Operators reported at 7:08 PM that the Klystron Debuncher cavity flow and LCW valve position were showing uncharacteristic oscillations; Operators contacted a Linac person who suggested cycling the cavity temperature regulation loop, which worked to resolve the oscillations. BRF15 tripped off at 9:03 PM; it reset. At 10:31 PM, BRF11 tripped off; it reset.

Sunday February 26

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 1:06 AM, KRF5 tripped off without an indication of a problem; after making several attempts to bring the station back online operators called in Linac personnel. The MI South Anode Power Supply tripped off at 6:52 AM; it reset.

During the day shift, at 12:54 PM, Linac personnel reported that there was still a serious problem with the KRF5 charging supply power supply.

During the evening shift, Linac personnel reported at 6:25 PM that the problem with KRF5 was in its charging supply transformer; they were going to replace the transformer with a spare Operators prepared MI for an access into MI-40 so that FESS personnel could finish their work on the MI-40 sump; they completed this job at 8:41 PM. KRF1 tripped off at 8:47 PM; it reset. L:QPS105 and 108 tripped off at 10:22 PM; they reset. At 10:58 PM, LRF2 tripped off; it reset. Linac personnel reported at 11:29 PM that they had completed the KRF5 charging supply power supply replacement.

Monday February 27

The midnight shift began with beam off due to KRF5 work.

During the midnight shift, Operations had beam established to NuMI and MiniBooNE by 1:34 AM. At 3:07 AM, Operations reported a few NuMI beam permit trips due to MI incomplete RF gap indications, but with no MI RF stations being off; Operators reset the trips. At 3:15 AM, the H- Source voltage dropped and had no output; Operators contacted Preaccelerator personnel, tried a few of their suggestions, and then asked them to come in. Preaccelerator personnel reported at 5:42 AM that they changed the power supply controller card. Operations had resumed sending beam to NuMI and MiniBooNE by 5:18 AM. At 5:57 AM, the MI-60 LCW system tank level alarmed on a low level indication. At 6:17 AM, BRF13 tripped off on a ground fault indication and BRF19 tripped off on RF gap faults; they both reset. At 7:26 AM, BRF16 tripped off; it reset. Operators reported at 7:52 AM that they had quite a bit of Booster RF sparking; they made some RF capture changes that minimized the sparking and station trips. At 7:57 AM, the Booster sparking resumed and caused BRF19 to trip off; it reset.

Operations Coordinator (OpsCo)

The OpsCo would like to hold off on a scheduled Booster RF station (BRF2) upgrade work until next week. Preaccelerator personnel need to access the H- Source, but that can wait until some naturally occurring downtime arises.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/27/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory